You bought a stock one year ago for $ 48.88 per share and sold it today for $ 56.03 per share. It paid a $ 1.44 per share dividend today. How much of the return came from dividend yield and how much came from capital​ gain?
Answer #1

Dividend yield=Dividend/Current value

=(1.44/48.88)=2.95%(Approx).

Capital gains yield=(End value-Beginning value)/Beginning value

=(56.03-48.88)/48.88=14.63%(Approx).